# Benchmarking quantum computers

### v1

@@ Abstract
The rapid pace of development in quantum computing technology has sparked a proliferation of benchmarks for assessing the performance of quantum computing hardware and software. Good benchmarks empower scientists, engineers, programmers, and users to understand a computing system’s power, but bad benchmarks can misdirect research and inhibit progress. In this Perspective, we survey the science of quantum computer benchmarking. We discuss the role of benchmarks and benchmarking, and how good benchmarks can drive and measure progress towards the long-term goal of useful quantum computations, i.e., “quantum utility”. We explain how different kinds of benchmark quantify the performance of different parts of a quantum computer, we survey existing benchmarks, critically discuss recent trends in benchmarking, and highlight important open research questions in this field.
In quantum computing, benchmarking means measuring the performance of quantum computing systems, and benchmarks are methods that are used to measure performance (Fig. 1a). The rapid advance of quantum computing hardware over the past decade has spawned a bewildering variety of concepts and techniques within the fledgling science of quantum computer benchmarking. In 2014, cutting edge “quantum computers” were physics experiments on a few qubits [1]. Today’s quantum computers routinely deploy 20-400 qubits [2, 3, 4, 5, 6] and can outperform classical computers on specially designed tasks [2]. But despite this progress, quantum computers are still in their infancy. Our field’s goal of solving problems of practical significance—often called “quantum utility”—remains elusive, and many scientists predict that achieving quantum utility will require years of progress and major technological advances [7, 8, 9, 10]. Benchmarking and benchmarks can measure and guide progress toward that goal.
@@
Benchmarks for classical computers, such as the LINPACK [11] and SPEC [12] benchmarks, measure proxies for computational utility. Once quantum computers become capable of computational utility, many quantum computer benchmarks will too. But today, and in the near term, quantum computer benchmarks should instead measure, guide, and incentivize progress towards utility (Fig. 1b). To solve practical problems that classical computers cannot, it is expected that quantum systems will need to grow to millions of physical qubits [10, 7, 9, 8]. But more qubits is not enough. Achieving quantum utility requires reducing and mitigating hardware errors, which are the imperfections in qubits and quantum logic gates that cause large quantum programs to fail (see Box 1). The importance of other limiting factors, like speed or power consumption, pales in comparison. Quantum logic operations fail far more often (today, typically around 0.01%-1% of the time [2, 3, 4, 5, 6]) than classical instructions (around $10^{-23}\%$ of the time [13]), and correcting or mitigating quantum errors is startlingly hard [14, 15, 16]. Because errors will dominate the performance and design of quantum computers for the foreseeable future, most quantum computer benchmarks measure the impacts of those errors.
@@
Many of the earliest and most widely-used quantum computer benchmarks, such as randomized benchmarking (RB) [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], directly measure the rates of errors in a quantum computer’s quantum logic gates. But quantum computers are not just arrays of qubits. Instead, a phalanx of powerful classical computing subcomponents—controllers, compilers, routers, and schedulers—manage and control the qubits [30]. Increasingly, quantum computations will be performed not directly on physical qubits, but instead on higher-performing logical qubits encoded within many physical qubits running a quantum error correction (QEC) algorithm [14]. These complexities have motivated an ever-increasing array of incomparable and complementary benchmarks that test different subsystems in a quantum computer and/or measure different performance metrics. Understanding and leveraging a benchmark requires knowing what systems it tests, what it measures about those subsystems’ performance, and why that metric of performance matters.
@@
In this perspective, we survey the science of quantum computer benchmarking and we highlight important open research questions in this field. Throughout, we limit our scope to the benchmarking of gate-model quantum computers built from qubits (the primary quantum computing paradigm), e.g., we do not discuss benchmarks relevant only to quantum annealers or analog quantum simulators.
@@
The field of quantum computer benchmarking evolved out of quantum tomography and error characterization [31, 32, 33]. These are techniques for reconstructing quantum states and learning detailed error models for qubits and quantum computers. The earliest “quantum computers” were physics experiments on one or two qubits, and physicists used characterization methods to learn about how their systems behaved [34, 35]. But tomography of $n$ qubits is exponentially expensive in $n$, so as $n$ grew it became necessary to summarize the performance of prototype quantum computers using a handful of metrics measured by benchmarks [17, 18].
A quantum computer benchmark (or just “benchmark” hereafter) is defined by (i) a set of computational tasks together with a procedure for performing those tasks on a quantum computer, and (ii) a method for computing one or more performance metrics from the (classical) data produced by running those tasks (Fig. 1a). This definition for a benchmark encompasses most uses of the term in quantum computer science, and it defines our scope. For example, benchmarks that directly measure materials properties (e.g., of an ion trap) are outside our scope.
For most benchmarks, the tasks comprise one or more quantum circuits, and for many benchmarks the output is a single, scalar metric of performance, e.g., quantum volume [36]. However, a single benchmark can compute multiple metrics (e.g., a success rate and a run time), or non-scalar quantities such as capability regions[37] (see Fig. 4). As with classical computers, no single benchmark or performance metric can capture all important aspects of a quantum computer’s performance.
Box 1. Errors in quantum computers. Gate-model quantum computers implement computations using quantum circuits, which are sequences of quantum logic gates [38] (see example below). When circuits are run on quantum computing hardware, errors can occur (red star, below). Errors cause quantum computations to fail by corrupting the outputs of quantum circuits, and most benchmarks quantify the magnitude and/or impact of these errors.[Uncaptioned image] A simple and popular model for errors in quantum circuits is that a random bit (and/or phase) flip occurs with some probability $\epsilon$ at each location in a circuit. So the probability of at least one error in a circuit of width $w$ and depth $d$ is $1-(1-\epsilon)^{wd}\lesssim wd\epsilon$, implying that a computation will likely succeed only if $\epsilon\ll 1/(wd)$. In this simple model, a quantum computer’s errors are described by a single error rate ($\epsilon$). But real quantum computers experience diverse and a priori unknown errors, and each kind of error has different effects [37, 39, 40]. Important examples of error types include coherent (i.e., systematic) and stochastic Pauli errors [39], crosstalk [41, 42], and slow drifts in a system’s parameters [43]. The complexities of the errors that occur in real quantum computers makes diverse, complementary benchmarks and metrics necessary.
@@
The field of quantum computer benchmarking began in the early 2000s with the initial development of randomized benchmarking (RB) [17, 18, 19] as well as pioneering experiments on 2-12 qubits using quantum algorithms for solving tiny problems as informal benchmarks [44, 45, 46, 47]. The purpose of all these benchmarks was to directly quantify the rates of errors associated with qubits and gates, and their impact on quantum circuits (Box 1 reviews errors and circuits). For example, RB measures the error rates of quantum gates. These earliest methods are low-level benchmarks (see Fig. 2) that isolate qubit and gate performance from classical components (e.g., compilers) within an integrated quantum computing system. Since 2005, many low-level benchmarks that measure gate and circuit error rates have been (and continue to be) developed. They complement these earliest methods or address their limitations, and we discuss several of them in the next section.
In the last decade, benchmarking research has increasingly focused on high-level benchmarks (see Fig. 2) that quantify the overall performance of integrated (a.k.a. full-stack) quantum computers. The most well-known high-level benchmark is IBM’s quantum volume benchmark [36]. High-level benchmarks still predominantly quantify the impact of errors in quantum gates and qubits, but they do so by measuring their effect within the context of a complete quantum computing “stack” (see Fig. 3) that potentially includes a lot of classical computing subsystems. High-level benchmarks can quantify the impact of errors from the perspective of a quantum computer user executing high-level quantum programs. But to do so, they mix together many different aspects of performance [30, 48], including the performance of purely classical subsystems, that may not all be relevant to the pursuit of quantum utility.
The trend towards high-level benchmarks and metrics over the last decade has been driven by hopes for near-term quantum utility with “NISQ” computing [49, 50, 51] (see Box 2), by the emergence of commercial quantum computing systems with tightly integrated classical software stacks [30], and by the needs of prospective quantum computer users with limited quantum computing expertise. In the last few years, this trend has culminated with the development of many benchmarking suites centered around quantum algorithms and applications [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 6, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. Although these trends are suggestive of a technology that is at or near utility, achieving utility seems like it will require major technological advances, notably ultra-reliable logical qubits (see Box 2). As long as quantum computing is in its infancy, high-level benchmarks—including those based on algorithms and applications—should be designed, used, and interpreted with caution.
@@
In principle, benchmarks are passive tools for measuring performance metrics, used to localize devices in “performance metric space” (see Fig. 1b). But benchmarks also exert sociological forces. They influence decisions both small (e.g., gate calibrations) and large (e.g., funding allocations), and drive resources towards optimizing performance on those benchmarks. Benchmarks are used to compare competing quantum computing technologies and make judgements (sometimes implicit) about the current and future merits of each technology [53]. They therefore influence the design of future quantum computing systems. Benchmarks are also used to optimize individual quantum computing systems, by choosing system settings (e.g., compilation parameters, routing algorithms, or gate pulses) that maximize performance on a benchmark [48]. This optimization can even be an automated part of regular calibration or tune-up routines [73, 74].
@@
Figure 2: Kinds of benchmark. Benchmarks vary widely in the abstraction level of their tasks, ranging from executing specific low-level quantum circuits to solving a computational problem, and by the complexity of the object whose performance they measure, ranging from individual logic gates to entire computing systems. The abstraction and complexity of a selection of important benchmarks and benchmark families are shown here.
@@
Using benchmarks to inform a quantum computing system’s design and settings can easily lead to bad choices, because it is unlikely that any small set of benchmarks can summarize all important aspects of performance. Fundamental design decisions primarily aimed at maximizing performance on any existing high-level benchmarks would be particularly shortsighted, because no existing high-level benchmarks have been shown to reliably quantify progress towards quantum utility. For example, quantum volume favors systems with high qubit connectivity [30, 48] but higher connectivity will not necessarily enable better logical qubits. Optimizing an existing system’s settings to maximize its performance on benchmarks is typically less consequential, but can also result in poor performance unless the benchmarks are matched to the intended uses of that system. For example, using (only) one- and two-qubit RB to calibrate the gates of a system intended to run many-qubit circuits (e.g., NISQ algorithms or QEC) is a bad choice if that system has significant crosstalk errors that could be reduced by calibrations, because these benchmarks are largely insensitive to these errors [75, 24].
@@
Widely adopted benchmarks and metrics that become de facto or official standards are particularly influential. Today, quantum volume is the most popular metric for summarizing an integrated quantum computer’s overall performance [76, 48]. A quantum computer’s total number of qubits and its average gate error rate (measured using RB) are the most popular metrics for summarizing its low-level performance. These three metrics are informative, but any small set of metrics has limitations. These metrics alone do not reliably quantify progress towards quantum utility (see later), implying that caution is necessary when using them to compare the quantum computing systems created by different companies and research groups.
The broad influence of benchmarks makes it critically important to design, adopt, and carefully use “good” benchmarks. Good benchmarks aid progress towards quantum utility (and eventually towards increasing utility) whereas bad benchmarks impede it. Benchmarks that measure technology-specific or highly technical metrics (e.g., leakage rates [77]) can and do aid this progress. But benchmarks that are technology independent and create easily digestible performance summaries (like quantum volume) will exert greater influence (for good or ill). Currently, the merits of those benchmarks should be primarily judged by whether they push the entire field of quantum computing towards computational utility.
Properties of good benchmarks
Good benchmarks are:
@@
    Well-defined. A benchmark should have an unambiguous procedure, i.e., any unspecified steps in that procedure should be intentional configurable parameters (e.g., some benchmarks allow creative compilation—see Box 4).
@@
These properties are not binary, and not every useful benchmark scores highly on all of them. For example, informal benchmarks are often created to demonstrate experimental advances and these benchmarks need only serve the scientific purposes at hand (so, e.g., implementation robustness is irrelevant). However, for a benchmark to warrant trust from a broad cross-platform community, it needs to satisfy the above properties. Otherwise, that benchmark may mislead and contribute to misdirection of resources. Few (and perhaps none) of today’s high-level benchmarks satisfy all the criteria listed above.
Benchmarks must be robust if they are to reliably compare different technologies, or positively influence engineering decisions. Benchmarks are ideally designed to measure an independently defined (and well-motivated) metric (e.g., gate fidelity), and a robust benchmark accurately measures the intended metric under broad conditions. However, proving that a benchmark is robust is often difficult, and this is the central task in much of the theory of benchmarks. The accepted standard is to prove that a benchmark accurately measures its metric for any system whose qubits experience small Markovian errors [39], as exemplified by the theory of RB [28, 78, 79, 80]. However, real physical qubits (and likely logical qubits) often experience significant non-Markovian errors, so theories that address more general errors (and perhaps reveal weaknesses in existing benchmarks) are a need for the field.
@@
Box 2. NISQ and fault-tolerant quantum computing. Quantum algorithms can be executed directly on physical qubits, which are not inherently fault tolerant. This approach, using systems akin to those available today, is referred to as noisy intermediate-scale quantum (NISQ) computing [49]. NISQ computers possess a moderate number of qubits, ranging from approximately 50 to 1000, and exhibit moderate error rates, between about $1\%$ and $0.1\%$. It is hoped that NISQ computers will demonstrate quantum utility through the use of heuristic algorithms specifically designed for such systems, commonly known as NISQ algorithms [49, 50, 51]. However, achieving utility-scale instances of quantum algorithms that exhibit a clear quantum advantage, such as Shor’s algorithm [81], is expected to require billions of gates [7, 8, 9, 10]. Consequently, executing these algorithms successfully necessitates error rates around $10^{-9}$ or lower, a target currently considered infeasible for physical qubits. Quantum computations can be rendered tolerant to errors by employing redundantly encoded logical qubits, protected using fault-tolerant quantum error correction (QEC) techniques [14]. According to fault tolerance theory, these logical qubits can achieve significantly lower error rates than their physical counterparts, provided the errors in the physical qubits are sufficiently rare and exhibit desirable characteristics (e.g., are adequately uncorrelated) [82]. In fault-tolerant quantum computing, physical qubits undergo repeated cycles of QEC, which differ significantly from the operations used in NISQ algorithms. This distinction necessitates different benchmarks and metrics for evaluating prototype NISQ and fault-tolerant quantum computers. Quantum algorithms are compiled very differently for NISQ and fault-tolerant architectures, and typically require many more gates (e.g., 50×50$\times$50 × more) for fault-tolerant execution. While most physical qubits can be directly manipulated using a continuous, universal set of one- and two-qubit gates [38], logical qubits are restricted to a discrete and non-universal set of “easy” gates, as per the Eastin-Knill theorem [15]. Achieving universal fault-tolerant quantum computation requires the use of “hard” gates, which require additional resources for implementation (e.g., magic state distillation [16]). This distinction has significant implications for the development of benchmarks for quantum computers.
@@
Many high-level benchmarks measure self-defined metrics (e.g., quantum volume is defined as the quantity measured by its eponymous benchmark [36]) and this complicates assessing those benchmarks’ robustness. Such benchmarks are always intended to measure proxies for some independent (if perhaps imprecisely defined) aspect of performance. For example, high-level benchmarks based on one or more algorithms are typically meant to quantify how well a system can run “interesting” instances of those algorithms, but they do not necessarily do so. The intended interpretation of a benchmark implies an imprecise notion of robustness that can be assessed. Arguably, the quantum volume benchmark is robust [30, 48], whereas many other existing high-level benchmarks are not. However, the inherent imprecision and subjectivity in assessing the merits of benchmarks with self-defined metrics is unfortunate. The field would benefit from the creation of high-level benchmarks that robustly measure well-motivated and independent performance metrics.
@@
Benchmark efficiency is necessary to make a benchmark practical, but it is surprisingly challenging to design efficient benchmarks. Many benchmarks are intended to be run on $n$ qubits for any user-specified $n$, and such a benchmark is formally efficient (or scalable) if it requires resources that grow as a polynomial in $n$. But, in practice, a stronger but less precise notion of efficiency (scalability) is often desirable: the resources required are small for all relevant $n$. The creation of efficient benchmarks is an increasing focus of the benchmarking research community, due to the ongoing rapid increase in the number of qubits in quantum computers. Scalable benchmarks include modern versions of RB [22, 23, 24, 25], cycle benchmarking [83], and mirror circuit benchmarks [37, 84]. However, there are many widely-used benchmarks (e.g., cross-entropy benchmarking [XEB] [85] and the quantum volume benchmark [36]) that are not scalable, because they rely on classical simulations of general quantum circuits (see Box 3).

### v2

@@ Abstract
The rapid pace of development in quantum computing technology has sparked a proliferation of benchmarks for assessing the performance of quantum computing hardware and software. Good benchmarks empower scientists, engineers, programmers, and users to understand a computing system’s power, but bad benchmarks can misdirect research and inhibit progress. In this Perspective, we survey the science of quantum computer benchmarking. We discuss the role of benchmarks and benchmarking, and how good benchmarks can drive and measure progress towards the long-term goal of useful quantum computations, i.e., “quantum utility”. We explain how different kinds of benchmark quantify the performance of different parts of a quantum computer, survey existing benchmarks, examine recent trends in benchmarking, and highlight important open research questions in this field.
In quantum computing, benchmarking means measuring the performance of quantum computing systems, and benchmarks are methods that are used to measure performance (Fig. 1a). The rapid advance of quantum computing hardware over the past decade has spawned a bewildering variety of concepts and techniques within the fledgling science of quantum computer benchmarking—a subfield of QCVV (quantum characterization, verification, and validation) [1]. In 2014, cutting edge “quantum computers” were physics experiments on a few qubits [2]. Today’s quantum computers routinely deploy 20-400 qubits [3, 4, 5, 6, 7] and can outperform classical computers on specially designed tasks [3]. But despite this progress, quantum computers are still in their infancy. Our field’s goal of solving problems of practical significance—often called “quantum utility”—remains elusive, and many scientists predict that achieving quantum utility will require years of progress and major technological advances [8, 9, 10, 11]. Benchmarking and benchmarks can measure and guide progress toward that goal.
@@
Benchmarks for classical computers, such as the LINPACK [12] and SPEC [13] benchmarks, measure proxies for computational utility. Once quantum computers become capable of computational utility, many quantum computer benchmarks will too. But today, and in the near term, quantum computer benchmarks should instead measure, guide, and incentivize progress towards utility (Fig. 1b). To solve practical problems that classical computers cannot, it is expected that quantum systems will need to grow to millions of physical qubits [11, 8, 10, 9]. But more qubits is not enough. Achieving quantum utility requires reducing and mitigating hardware errors, which are the imperfections in qubits and quantum logic gates (simply “gates” hereafter) that cause large quantum programs to fail (see Box 1). The importance of other limiting factors, like speed or power consumption, pales in comparison. Quantum logic operations fail far more often (today, typically around 0.01%-1% of the time [3, 4, 5, 6, 7]) than classical instructions (around $10^{-23}\%$ of the time [14]), and correcting or mitigating quantum errors is startlingly hard [15, 16, 17, 4, 18, 19, 20]. Because errors will dominate the performance and design of quantum computers for the foreseeable future, most quantum computer benchmarks measure the impacts of those errors.
@@
Many of the earliest and most widely-used quantum computer benchmarks, such as randomized benchmarking (RB) [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], directly measure the rates of errors in a quantum computer’s gates. But quantum computers are not just arrays of qubits. Instead, a phalanx of powerful classical computing subcomponents—controllers, compilers, routers, and schedulers—manage and control the qubits [34]. Increasingly, quantum computations will be performed not directly on physical qubits, but instead on higher-performing logical qubits encoded within many physical qubits running a quantum error correction (QEC) algorithm [15]. These complexities have motivated an ever-increasing array of incomparable and complementary benchmarks that test different subsystems in a quantum computer and/or measure different performance metrics. Understanding and leveraging a benchmark requires knowing what systems it tests, what it measures about those subsystems’ performance, and why that metric of performance matters.
@@
In this perspective, we survey the science of quantum computer benchmarking and we highlight important open research questions in this field. Our focus is on benchmarking rather than the broader field of QCVV, and we present a forward-looking and opinionated perspective that complements recent reviews and tutorials on QCVV [1, 35, 36] and Carrasco et al.’s perspective on verifying results from quantum devices [37]. Throughout, we limit our scope to the benchmarking of gate-model quantum computers built from qubits (the primary quantum computing paradigm), e.g., we do not discuss benchmarks relevant only to quantum annealers or analog quantum simulators. We presume that the goal of quantum computing R&D is to achieve and then advance computational utility, and we measure the relative value of benchmarks by how useful and appropriate they are to that goal.
We present our perspective in three sections. In the first section, we discuss the origins of quantum computer benchmarking, current trends in the field, and the roles that benchmarks play within the quantum computing community. Having summarized the current state of play, we then discuss what properties make a benchmark broadly useful. In the second section, we provide a framework within which benchmarks can be categorized. We use this framework to analyze many of the most important and popular benchmarks and highlight specific gaps in the benchmarking toolbox. In the final section, we turn to what is, in our view, the largest open question in benchmarking: how can our field create and deploy benchmarks that reliably quantify progress towards quantum utility? We discuss how these benchmarks might be created, presenting our view on the ingredients needed from across the quantum computing community.
Box 1. Errors in quantum computers. Gate-model quantum computers implement computations using quantum circuits, which are sequences of gates [38] [see example below, containing the Hadamard (H), phase (P) and CNOT gates (connected circles)]. When circuits are run on quantum computing hardware, errors can occur (red star, below). Errors cause quantum computations to fail by corrupting the outputs of quantum circuits, and most benchmarks quantify the magnitude and/or impact of these errors.[Uncaptioned image] A simple and popular model for errors in quantum circuits is that a random bit (and/or phase) flip occurs with some probability $\epsilon$ at each location in a circuit. So the probability of at least one error in a circuit of width $w$ and depth $d$ is $1-(1-\epsilon)^{wd}\lesssim wd\epsilon$, implying that a computation will likely succeed only if $\epsilon\ll 1/(wd)$. In this simple model, a quantum computer’s errors are described by a single error rate ($\epsilon$). But real quantum computers experience diverse and a priori unknown errors, and each kind of error has different effects [39, 40, 41]. Important examples of error types include coherent (i.e., systematic) and stochastic Pauli errors [40], crosstalk [42, 43], and slow drifts in a system’s parameters [44]. The complexities of the errors that occur in real quantum computers makes diverse, complementary benchmarks and metrics necessary.
@@
The field of quantum computer benchmarking evolved out of quantum tomography and error characterization [45, 46, 47, 48, 49]. These are techniques for reconstructing quantum states and learning detailed error models for qubits and quantum computers. The earliest “quantum computers” were physics experiments on one or two qubits, and physicists used characterization methods to learn about how their systems behaved [50, 51]. Early benchmarks (e.g., RB [21, 23, 22]) emerged to provide simpler, more concise assessments of one or two qubit operations. As $n$-qubit systems with $n\geq 5$ emerged, full tomographic characterization of all $n$ qubits grew rapidly infeasible. It became both convenient and necessary to summarize the “holistic” performance of prototype quantum computers using a handful of metrics measured by benchmarks [52, 53, 54, 55, 56, 25].
A quantum computer benchmark (or just “benchmark” hereafter) is defined by (i) a set of computational tasks together with a procedure for performing those tasks on a quantum computer (or one or more of its components), and (ii) a method for computing one or more performance metrics from the (classical) data produced by running those tasks (Fig. 1a). This definition for a benchmark encompasses most (but not all) uses of the term in quantum computer science, and it defines our scope. Benchmarks that measure metrics that are not direct computational properties of quantum computing systems (e.g., qubit coherence times, or heating rates of ion traps) are outside our scope, including benchmarks that measure a qubit’s $T_{1}$ and $T_{2}$ times [1], which are widely-used metrics for qubit quality.
For most benchmarks, the tasks comprise one or more quantum circuits, and for many benchmarks the output is a single, scalar metric of performance, e.g., quantum volume [52]. However, a single benchmark can compute multiple metrics (e.g., a success rate and a run time), or non-scalar quantities such as capability regions[39] (see Fig. 4). In our view, no single benchmark or performance metric can capture all important aspects of a quantum computer’s performance, as is accepted with classical computers.
@@
The field of quantum computer benchmarking began in the early 2000s with the initial development of RB [21, 22, 23] as well as pioneering experiments on 2-12 qubits using quantum algorithms for solving tiny problems or creating canonical states as informal benchmarks [57, 58, 59, 60, 61]. The purpose of all these benchmarks was to directly quantify the rates of errors associated with qubits and gates, and their impact on quantum circuits (Box 1 reviews errors and circuits). For example, RB measures the error rates of gates. These earliest methods are low-level benchmarks (see Fig. 2) that isolate qubit and gate performance from classical components (e.g., compilers) within an integrated quantum computing system. Since 2005, many low-level benchmarks that measure gate and circuit error rates have been (and continue to be) developed. They complement these earliest methods or address their limitations, and we discuss several of them in the next section.
In the last decade, benchmarking research has increasingly focused on high-level benchmarks (see Fig. 2) that quantify the overall performance of integrated (a.k.a. full-stack) quantum computers. The most well-known high-level benchmark is IBM’s quantum volume benchmark [52]. High-level benchmarks still predominantly quantify the impact of errors in gates and qubits, but they do so by measuring their effect within the context of a complete quantum computing “stack” (see Fig. 3) that potentially includes a lot of classical computing subsystems. High-level benchmarks can quantify the impact of errors from the perspective of a quantum computer user executing high-level quantum programs. But to do so, they mix together many different aspects of performance [34, 62], including the performance of purely classical subsystems, that may not all be relevant to the pursuit of quantum utility.
The trend towards high-level benchmarks and metrics over the last decade has been driven by hopes for near-term quantum utility with noisy intermediate-scale quantum (NISQ) computing [63, 64, 65] (see Box 2), by the emergence of commercial quantum computing systems with tightly integrated classical software stacks [34], and by the needs of prospective quantum computer users with limited quantum computing expertise. In the last few years, this trend has culminated with the development of many benchmarking suites centered around quantum algorithms and applications [66, 53, 67, 68, 69, 70, 71, 72, 73, 74, 75, 7, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85]. Although these trends are suggestive of a technology that is at or near utility, achieving utility seems like it will require major technological advances, notably ultra-reliable logical qubits (see Box 2). As long as quantum computing is in its infancy, our opinion is that high-level benchmarks—including those based on algorithms and applications—should be designed, used, and interpreted with caution.
@@
In principle, benchmarks are passive tools for measuring performance metrics, used to localize devices in “performance metric space” (see Fig. 1b). But benchmarks also exert sociological forces. They influence decisions both small (e.g., gate calibrations) and large (e.g., funding allocations), and drive resources towards optimizing performance on those benchmarks. Benchmarks are used to compare competing quantum computing technologies and make judgements (sometimes implicit) about the current and future merits of each technology [53, 6, 73, 68, 81, 80, 70]. They therefore influence the design of future quantum computing systems. Benchmarks are also used to optimize individual quantum computing systems, by choosing system settings (e.g., compilation parameters or gate pulses) that maximize performance on a benchmark [62, 86, 87], sometimes with automated routines [86, 87, 88].
@@
Figure 2: Kinds of benchmark. Benchmarks vary widely in the abstraction level of their tasks, ranging from executing specific low-level quantum circuits to solving a computational problem, and by the complexity of the object whose performance they measure, ranging from individual gates to entire computing systems. The abstraction and complexity of a selection of important benchmarks and benchmark families are shown here. We also show tomographic (characterization) methods (e.g., state, process, and gate set tomography), which are not benchmarks, for comparison.
@@
Using benchmarks to inform a quantum computing system’s design and settings can easily lead to unproductive choices, because the chosen benchmarks might not summarize all important aspects of performance. Fundamental design decisions primarily aimed at maximizing performance on any existing high-level benchmarks would be particularly shortsighted in our view, because no existing high-level benchmarks have been shown to quantify progress towards quantum utility. For example, quantum volume favors systems with high qubit connectivity [34, 62, 89, 6] but higher connectivity will not necessarily enable better logical qubits. Optimizing an existing system’s settings to maximize its performance on benchmarks is less consequential, but can also result in poor performance unless the benchmarks are matched to that system’s intended uses. For example, using (only) one- and two-qubit RB to calibrate the gates of a system [86, 87, 88] intended to run many-qubit circuits (e.g., NISQ algorithms or QEC) is unhelpful if that system has significant crosstalk errors [42, 43, 28, 90] that could be reduced by calibrations, because these benchmarks are largely insensitive to those errors [91, 28].
@@
Widely adopted benchmarks and metrics that become de facto or official standards are particularly influential. Today, quantum volume is the most popular metric for summarizing an integrated quantum computer’s overall performance [89, 62]. A quantum computer’s total number of qubits and its average gate error rate (measured using RB) are the most popular metrics for summarizing its low-level performance. These metrics are informative, but any small set of metrics has limitations. These metrics alone do not reliably quantify progress towards quantum utility (see later), so our opinion is that caution is needed when using them to compare quantum computing systems.
Desirable properties of benchmarks
The broad influence of benchmarks makes it critically important to design, adopt, and carefully use “good” benchmarks. In our view, a benchmark is good if it aids progress towards quantum computational utility (and eventually towards increasing utility) when used correctly, and if it is difficult to misuse. We do not know any algorithm for guaranteeing that a benchmark will satisfy these desiderata. But we can identify desirable properties that, in our opinion and experience, tend to make benchmarks useful and discourage misuse. Benchmarks that do not satisfy these properties are less likely to aid progress toward utility, and are more easy to misuse and misinterpret.
@@
    Well-defined. A benchmark should have an unambiguous procedure, i.e., any unspecified steps in that procedure should be intentional configurable parameters (e.g., some benchmarks allow creative compilation—see Box 3).
@@
Box 2. NISQ and fault-tolerant quantum computing. Quantum algorithms can be executed directly on physical qubits, which are not inherently fault tolerant. This approach, using systems akin to those available today, is referred to as noisy intermediate-scale quantum (NISQ) computing [63]. NISQ computers possess a moderate number of qubits, ranging from approximately 20 to 1000, and exhibit moderate error rates, between about $1\%$ and $0.1\%$. It is hoped that NISQ computers will demonstrate quantum utility through the use of heuristic algorithms specifically designed for such systems, commonly known as NISQ algorithms [63, 64, 65]. However, achieving utility-scale instances of quantum algorithms that exhibit a clear quantum advantage, such as Shor’s algorithm [92], is expected to require billions of gates [8, 9, 10, 11]. Consequently, executing these algorithms successfully necessitates error rates around $10^{-9}$ or lower, a target currently considered infeasible for physical qubits. Quantum computations can be rendered tolerant to errors by employing redundantly encoded logical qubits, protected using fault-tolerant quantum error correction (QEC) techniques [15, 4, 18, 19, 20]. According to fault tolerance theory, these logical qubits can achieve significantly lower error rates than their physical counterparts, provided the errors in the physical qubits are sufficiently rare and exhibit desirable characteristics (e.g., are adequately uncorrelated) [93]. In fault-tolerant quantum computing, physical qubits undergo repeated cycles of QEC, which differ significantly from the operations used in NISQ algorithms. This distinction necessitates different benchmarks and metrics for evaluating prototype NISQ and fault-tolerant quantum computers. Quantum algorithms are compiled very differently for NISQ and fault-tolerant architectures, and typically require many more gates (e.g., $50\times$ more) for fault-tolerant execution. While most physical qubits can be directly manipulated using a continuous, universal set of one- and two-qubit gates [38], logical qubits are restricted to a discrete and non-universal set of “easy” gates, as per the Eastin-Knill theorem [16]. Achieving universal fault-tolerant quantum computation requires the use of “hard” gates, which require additional resources for implementation (e.g., magic state distillation [17]). This distinction has significant implications for the development of benchmarks for quantum computers.
These properties describe ideals against which we think it is useful to judge benchmarks. But they are not precise or binary properties, and not every useful benchmark scores highly on all of them. For example, informal benchmarks are often created to demonstrate experimental advances and these benchmarks need only serve the scientific purposes at hand (so, e.g., implementation robustness is irrelevant). However, a benchmark warrants trust from the entire quantum computing community only if it approximates these ideals. In our opinion, few (and probably none) of today’s high-level benchmarks achieve this standard.
Benchmarks must be implementation- and system-robust if they are to enable reliable comparisons between different technologies, or positively influence engineering decisions. Benchmarks are ideally designed to measure independently defined and well-motivated metrics (e.g., gate fidelity), and such benchmarks are robust if they accurately measure the intended metrics under broad conditions. However, proving that a benchmark is robust is often difficult, and is the central task in much of the theory of benchmarks. The accepted standard is to prove that a benchmark accurately measures its metric for any system whose qubits experience small Markovian errors [40], as exemplified by the theory of RB [32, 94, 95, 96]. However, qubits often experience significant non-Markovian errors, and those errors can often corrupt existing benchmarks (e.g., leakage can corrupt RB [97]). Theories and benchmarks that address more general errors are beginning to be developed [98, 99, 97, 44] but remain a need for the field.
@@
Many high-level benchmarks measure self-defined metrics (e.g., quantum volume is defined as the quantity measured by its eponymous benchmark [52]) and this complicates assessing those benchmarks’ robustness. Such benchmarks are always intended to measure proxies for some independent (if perhaps imprecisely defined) aspect of performance. For example, high-level benchmarks based on one or more algorithms are typically meant to quantify how well a system can run “interesting” instances of those algorithms (or how far the system is from running those interesting instances), but they do not necessarily do so. The intended interpretation of a benchmark implies an imprecise notion of robustness that can be assessed. Arguably, the quantum volume benchmark is robust [34, 62], whereas many other existing high-level benchmarks are not. However, the inherent imprecision and subjectivity in assessing the merits of benchmarks with self-defined metrics is unfortunate. The field would benefit from the creation of high-level benchmarks that robustly measure well-motivated and independent performance metrics.
Benchmark efficiency is necessary to make a benchmark practical, but it is surprisingly challenging to design efficient benchmarks. Many benchmarks are intended to be run on $n$ qubits for any user-specified $n$, and such a benchmark is formally efficient (or scalable) if it requires resources that grow as a polynomial in $n$. But, in practice, a less precise notion of efficiency (scalability) is often more important: the resources required are small for all relevant $n$. The creation of efficient benchmarks is an increasing focus of the benchmarking research community, due to the ongoing rapid increase in the number of qubits in quantum computers. Scalable benchmarks include modern versions of RB [26, 27, 28, 29], cycle benchmarking [56], and mirror circuit benchmarks [39, 100]. However, there are many widely-used benchmarks (e.g., cross-entropy benchmarking [XEB] [55] and the quantum volume benchmark [52]) that are not scalable, because they rely on classical simulations of general quantum circuits (see Box 4).